OpenStudy (anonymous):

if h(x) = x^2 + 5, then what is h^-1(x) (the inverse)?

OpenStudy (anonymous):

h=x^2+5 switch variable x=h^2+5 x-5=h^2 h=+ or -(x-5)

OpenStudy (anonymous):

h=+ or -Sqrt[(x-5)]

OpenStudy (anonymous):

thank you!